OpenStudy (anonymous):

A rectangle has a length of 12 and width 6. How long is the diagonal?

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

\(\bf c^2=a^2+b^2\implies c=\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

 But I'm still confused because my teacher said to do it this way and I keep getting the wrong answer... I keep getting 12 but that doesn't make sense.